40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 5961     Entry time: Sat Nov 19 15:58:04 2011
Author: Mirko 
Type: Update 
Category: IOO 
Subject: Some more looks into OSEM noise 

[Den, Mirko]

We looked some more into the the OSEM signals and their coherence to the seismometer signals.

We were able to verify that the coherence OSEM sensor <-> seismometer signal goes down with increasing the OSEM gain. This seems to indicate that the OSEM FB add noise to the distance mirror <-> frame. We injected some noise into the OSEMs to see how the coherence behaves.

MC2 SUSPOS, 0.1Hz - 0.8Hz, 3mins each

Inj. amplitude   Time(UTC) Note

-                     21:35          Free swinging
-                     21:42          Big LF OSEM gain
-                     21:48          Small LF OSEM gain
150                     21:56          -"-
300                     22:00          -"-
900                     22:05          -"-

Free swinging:


High OSEM gain:


Low OSEM gain:






We left the filters that lower the OSEM gain below 0.3Hz on.

Attachment 2: High_Osem_Gain.pdf  19 kB  | Hide | Hide all
Attachment 4: Low_LF_OSEM_Gain.fig  306 kB
ELOG V3.1.3-